For each of the following solutions, how many milliliters of water should be added to...


8-67 For each of the following solutions, how many milliliters of water should be added to yield a solution that has a concentration of 0.100 M? 


8-69 Determine the final concentration of each of the following solutions after 20.0 mL of water has been added. 

a. 30.0 mL of 5.0 M NaCl solution b. 30.0 mL of 5.0 M AgNO, solution

c. 30.0 mL of 7.5 M NaCl solution d. 60.0 mL of 2.0 M NaCl solution 


8-115 At a given temperature, which of the two solutions 1.0 M glucose or 1.0 M NaCl has the higher 

a. vapor pressure b. boiling point c. freezing point d. osmotic pressure  
Answer #1

8.67)

A) M1V1 = M2V2

50×3 = 0.1×V2

V2 = 1500 mL

Water added = 1500-50 = 1450 mL

8.115)

Vapour pressure = Glucose

Boiling point = NaCl

Freezing point = glucose

Osmotic pressure = NaCl